write the slope-intercept equation for the line that passes through (-5,-8) and is perendicular to 10x-6y=-11 please show all of your work
Answer by mananth(16949)

10x-6y= -11
Find the slope of this line

-6y=-10 x-11
Divide by -6
y = 5/ 3 x + 11/6
Compare this equation with y=mx+b
slope m =5/3

The slope of a line perpendicular to the above line will be the negative reciprocal -3/5 m1*m2=-1


m= - 3/5 ,point ( -5 , -8 )
Find b by plugging the values of m & the point in
y=mx+b
-8 = 3.00 + b
b= -11
m= - 3/5
The required equation is y =- ( 3/ 5)x-11
m.ananth@hotmail.ca